# Detect Everything with Few Examples

### v3

@@ 4 Experiments
We comprehensively evaluate our method on few-shot and one-shot benchmarks. Furthermore, we compare the efficiency of our method against SoTA solutions, study few-shot performance for different numbers of shots, compare it to language-based detectors, and provide qualitative results. We conduct ablations to study the contributions of the proposed components to the performance of our method. Our source code and the pretrained models will be publicly released upon acceptance.
Evaluation Metrics and Datasets. Few-shot and one-shot evaluations split classes into base and novel classes. Base classes are seen during training and novel classes are unseen. The performance on novel classes is more important. For COCO and Pascal VOC, nAP, nAP50, and nAP75 represent mAP, AP50, and AP75 in novel classes. bAP and bAP50 represent mAP and AP50 in base classes. One-shot evaluation conventionally divides 80 classes of COCO into four even partitions, and alternatively uses three as base classes and one partition as novel classes [37]. There are 4 base/novel splits in total, named Split-1/2/3/4. For LVIS, APr, APc, and APf represent AP on rare, common, and frequent categories. Rare categories are used as novel classes. Metrics on LVIS, such as box APr and mask APr, are computed on bounding boxes and on instance segmentation masks separately. We evaluate our method on Pascal VOC [9], COCO [31], and LVIS-v1 [15]. We follow the conventional base/novel classes split and use the same support images of novel classes with existing work [46, 52].
@@
Figure 6: Qualitative visualization of our method DE-ViT on COCO (a) and LVIS (b), and comparison to Meta-Faster RCNN [18]. Boxes of base and novel classes are colored in green and yellow, respectively.
Implementation Details. We adopt a standard two-stage object detection framework, similarly to Mask R-CNN [20]. We use the same off-the-shelf RPN with existing work [56] to generate object proposals. The RPN is trained separately for each dataset using only base classes. We use DINOv2 [39] ViT as the backbone unless specified, and report results in ViT-S/B/L (small, base, large) model sizes. The ViT backbones are kept frozen during detector training. We adopt the prototype extraction procedure of Meta RCNN [51] using ViT features. Prototypes of background classes set $\mathcal{B}$ are extracted from images of stuff (non-object) classes, e.g., sky, road, from COCOStuff [3] unless specified. Similar to [51], the top $T$ most likely classes for each proposal are determined by the distances between prototypes and the average proposal feature. We set $T$ to 10 unless specified, where $T$ is the number of classes to create subspace features and perform inference as explained in Sec. 3.4. We apply 3 PL blocks for experiments on Pascal VOC and COCO, and 5 PL blocks for those on LVIS.
@@ 4.1 Main Results
Tab. 2 shows our results on the few-shot Pascal VOC benchmark. Our method DE-ViT outperforms the previous SoTA NIFF [14] by +2.0 on averaged nAP50. It is worth noting that the implementation of NIFF has not been released.
LVIS has been regarded as a highly challenging benchmark for FSOD [46] with 337 novel classes and only DiGeo [34] reports few-shot results on LVIS v1. Tab. LABEL:tab:lvis-fewshot shows that our method outperforms DiGeo in all metrics and a significant boost in the accuracy of detecting novel objects (+20 box APr).
Tab. 4 shows our results on the one-shot COCO benchmark. DE-ViT surpasses the previous SoTA BHRL by 6 bAP50 and 2.8 nAP50. One-shot object detection task follows a single-class detection setting. Therefore, we adapt our method DE-ViT by detecting each class separately during evaluation. OWL-ViT [38] also reports one-shot results on COCO. However, OWL-ViT’s results are obtained with an ensemble of language-based detection and one-shot pipelines without providing an implementation or isolated measurements. Therefore, we exclude OWL-ViT from the one-shot comparison.
@@
4.2 Analysis and Ablation Studies
@@
Efficiency. We compare the inference time of DE-ViT with different values of $T$ against recent few-shot works in Tab. 5 in COCO 10-shots setting, where Swin denotes Swin Transformer [32], RN101 denotes ResNet101 [21], MFRCNN denotes Meta Faster RCNN [18], and CrossT denotes Cross Transformer [19]. $T$ is the number of classes to create subspace features and perform inference. Tab. 5 shows that our method DE-ViT can achieve the smallest inference time while having better accuracy. The inference times of all the compared methods are measured on the same machine.
@@
Table 5: Ablation study on different values of $T$ and inference time comparison with existing methods. N/A: The pretrained model is not publicly available for evaluation.
@@
Qualitative results. Fig. 4 shows examples of region propagations on classes “horse” and “bird”. The regions gradually move from the initial proposal region toward the object regions and are transformed into predicted boxes with our spatial integral layer. Fig. 6 shows a qualitative demonstration of our method on COCO and LVIS, and a comparison to Meta Faster RCNN [18]. Note that we chose Meta Faster RCNN over other existing works because of the availability of its pretrained models. DE-ViT detects more novel objects while having fewer false positives. We provide more visualizations in the supplementary material.
Table 6: Comparison to language-based detectors on LVIS and COCO 2017. $\dagger$: use customized pretrained backbone instead of public ones. “-”: result that was not reported.
@@
DE-ViT (Ours, 30-shots) ViT-S/14 24.2 23.4 39.5
@@
Comparison to language-based detectors. Tab. 6 compares our method DE-ViT against language-based detectors on the COCO and LVIS-v1 datasets. Language-based detection, also known as open-vocabulary detection [48], does not assume access to support images on novel classes but requires the knowledge of the novel class names. The class names are then used to discover images of novel classes from a large collection of image-text pairs. Therefore, language-based detectors have the following limitations. First, many objects lack clear names [30], e.g., objects that are specific to certain contexts [24]. Second, the association between visual concepts and language is evolving and not static [8]. In contrast, few-shot object detection (FSOD) does not assume any knowledge of class names, and novel classes are described with support images only. FSOD’s setting is therefore more general and arguably more challenging, and aims to emulate humans’ capability to recognize objects by their consistent appearance.
@@
In COCO, our method DE-ViT outperforms the previous SoTA CORA+ by 6.9 AP50. Our method only trains on COCO while CORA+ uses ImageNet-21K [29] and COCO Captions [7] as additional training data. In LVIS, DE-ViT outperforms the previous SoTA on mask APr (+1.5 over F-VLM) and box APr (+2.4 over OWL-ViT). Note that we report the Co-Det [33] result with Swin-B backbone instead of EVA02-L backbone [12] because the latter includes the base and novel classes of COCO in its training set. Moreover, we observe a high variance in the performance of language-based detectors. F-VLM achieves a mask APr of 32.8 on LVIS but only has 28 nAP50 on COCO. While CORA+ has 43.1 nAP50 on COCO but only a box APr of 28.1 on LVIS. On the contrary, our method DE-ViT outperforms existing solutions on both LVIS and COCO.
Table 7: Ablation studies on the feature subspace projection.
Table 9: Ablation studies on the region propagation-based localization.
o c|ccc|ccccRaw Feature Input Feature Subspace Projection Novel Base
Prototype Classifier       $c$          $\mathcal{B}$    $\mathcal{C}\setminus c$ nAP50 nAP75 bAP50 bAP75
4.5 2.2 48.9 22.5
      ✓ 26.2 9.7 29.3 12
      ✓          ✓ 38.4 23 43.4 26.8
      ✓          ✓    ✓ 39.5 24.1 42.3 25.9
Conventional Expanded Region Propagation Novel Base
Regression Net Proposal Network nAP50 nAP75 bAP50 bAP75
37.7 14.6 46.5 23.9
35.6 12.5 41.3 19.8
39.5 24.1 42.3 25.9
Feature subspace projection. We examine the component effects for feature subspace projection in Tab. 4.2. The first column represents the integration of raw features of DINOv2 ViT into Meta RCNN [51], a standard prototype-based few-shot detector. Tab. 4.2 shows that the model has an accuracy collapse on novel classes with raw ViT feature inputs by completely overfitting base classes. The $c$, $\mathcal{B}$, and $\mathcal{C}\setminus c$ represents different groups of prototypes used in feature projection, as in Eq. 5. By projecting features to $p_{c}$, general detection ability emerges (nAP50: 4.5 $\rightarrow$ 26.2), then significantly improves after adding background classes prototypes $p_{\mathcal{B}}$ (nAP50: 26.2 $\rightarrow$ 38.4), and further improves with $p_{\mathcal{C}\setminus c}$ (nAP50: 38.4 $\rightarrow$ 39.5). Results in Tab. 4.2 and 4.2 are obtained on COCO 2017 with ViT-S/14.
Region propagation network. We study the impacts of region-propagation-based localization in Tab. 4.2. The conventional regression network exhibits poor localization accuracy on novel classes (nAP75: 14.6), and simply expanding proposals even lowers the performance (nAP75: 14.6 $\rightarrow$ 12.5). When using our proposed region-propagation network with the learnable spatial integral layer, the localization accuracy has a significant boost (nAP75: 12.5 $\rightarrow$ 24.1). Note that the results in Tab. 4.2 are obtained by only changing the localization architectures while keeping the feature subspace projection.
@@
Background classes set $\mathcal{B}$. We study the impacts of using different background classes sets $\mathcal{B}$ in Tab. LABEL:tab:abla-B. The first and second rows represent using the stuff (non-object) classes, e.g., floor, sky, of COCOstuff [3] and ADE20k [57] to construct $\mathcal{B}$. Note that COCOstuff has 53 stuff classes and ADE20k has 35 stuff classes, which explains the minor accuracy drop from ADE20k. The third row denotes using all classes of ADE20k as $\mathcal{B}$. This indicates that adding thing (object) classes to $\mathcal{B}$ may not be beneficial. Tab. LABEL:tab:abla-B shows that changing $\mathcal{B}$ only results in small accuracy variations. This suggests that our proposed feature subspace projection does not rely on specific prototypes.
Supervised object detection through region propagation. We further study the application of our proposed techniques in supervised object detection. In Fig. LABEL:fig:reg-loss, we compare the mAPs and regression loss curves of a standard RCNN head and using region propagation for RCNN prediction in COCO detection. We use a ResNet-50 backbone and the same number of parameters for each head. Despite having similar mAPs (37.15 vs. 37.23), Fig. LABEL:fig:reg-loss shows that our proposed region propagation produces a consistently lower regression loss than the standard RCNN. This suggests the potential of our proposed techniques in object detection tasks.
We provide further analysis on the effects of the number of shots, scenarios when expanded proposals cover multiple same-class objects, and more qualitative visualizations in the supplementary document.
Table 9: Ablation studies on the region propagation-based localization.

### v4

@@ 3 Experiments
We comprehensively evaluate our method on few-shot and one-shot benchmarks. Furthermore, we compare the efficiency of our method against SoTA solutions, study few-shot performance for different numbers of shots, compare it to language-based detectors, and provide qualitative results. We conduct ablations to study the contributions of the proposed components to the performance of our method. Our source code and the pretrained models are included in the supplementary material and will be publicly released upon acceptance.
Evaluation Metrics and Datasets. Few-shot and one-shot evaluations split classes into base and novel classes. Base classes are seen during training and novel classes are unseen. The performance on novel classes is more important. For COCO and Pascal VOC, nAP, nAP50, and nAP75 represent mAP, AP50, and AP75 in novel classes. bAP and bAP50 represent mAP and AP50 in base classes. One-shot evaluation conventionally divides 80 classes of COCO into four even partitions, and alternatively uses three as base classes and one partition as novel classes [30]. There are 4 base/novel splits in total, named Split-1/2/3/4. For LVIS, APr, APc, and APf represent AP on rare, common, and frequent categories. Rare categories are used as novel classes. Metrics on LVIS, such as box APr and mask APr, are computed on bounding boxes and on instance segmentation masks separately. We evaluate our method on Pascal VOC [7], COCO [8], and LVIS-v1 [9]. We follow the conventional base/novel classes split and use the same support images of novel classes [13, 11].
@@
Figure 6: Qualitative visualization of our method DE-ViT on COCO (a) and LVIS (b), and comparison to Meta-Faster RCNN [15]. Boxes of base and novel classes are colored in green and yellow.
Implementation Details. We adopt a standard two-stage object detection framework, similarly to Mask R-CNN [31]. We use the same off-the-shelf RPN with existing work [32] to generate object proposals. The RPN is trained separately for each dataset using only base classes. We use DINOv2 [33] ViT as the backbone, and report results in ViT-S/B/L (small, base, large) model sizes. The ViT backbones are kept frozen during detector training. We adopt the prototype extraction procedure of Meta RCNN [17] using ViT features. During training, the model can access only the prototypes of base classes. After training, the prototypes of novel classes are appended while the remaining parameters are unchanged. During evaluation, the model is evaluated on images that contain objects of both base and novel classes. Prototypes of background classes set $\mathcal{B}$ are extracted from images of stuff (non-object) classes, e.g., sky, road, from COCOStuff [34] unless specified. Similar to [17], the top $T$ most likely classes for each proposal are determined by the distances between prototypes and the average proposal feature. We set $T$ to 10 unless specified, where $T$ is the number of classes to create subspace features and perform inference as explained in Sec. 2.4. We apply 3 PL blocks for experiments on Pascal VOC and COCO, and 5 PL blocks for those on LVIS.
@@ 3.1 Main Results
Tab. 2 shows our results on the few-shot Pascal VOC benchmark. Our method DE-ViT outperforms the previous SoTA NIFF [12] by +2.0 on averaged nAP50. It is worth noting that the implementation of NIFF has not been released. LVIS has been regarded as a highly challenging benchmark for FSOD [13] with 337 novel classes and only DiGeo [14] reports few-shot results on LVIS v1. Tab. LABEL:tab:lvis-fewshot shows that our method outperforms DiGeo in all metrics and a significant boost in the accuracy of detecting novel objects (+20 box APr).
Tab. 3.1 shows our results on the one-shot COCO benchmark. DE-ViT surpasses the previous SoTA BHRL by 6 bAP50 and 2.8 nAP50. One-shot detection task follows a single-class detection setting. Therefore, we adapt our method DE-ViT by detecting each class separately during evaluation. OWL-ViT [36] also reports one-shot results on COCO. However, OWL-ViT’s results are obtained with an ensemble of language-based detection and one-shot pipelines without providing an implementation or isolated measurements. Therefore, we exclude OWL-ViT from the one-shot comparison.
@@
3.2 Real Robot Experiment
Figure 8: Overview of the system for sorting novel objects built upon our DE-ViT. Our system receives a few example images of novel classes and instantly detects new objects within the same category without tuning or further learning.
Table 5: Success rates of DE-ViT on sorting novel objects.
Object Success
Overall (%) 97%
Chips 8/10
Tomato Can 9/10
Ball 10/10
Crackers
Brick
Cup
Mustard Bottle
Sugar Box
Orange
Cleanser Bottle
Setup. To evaluate our DE-ViT in real-world robotic tasks, we develop a pick-and-place system for sorting novel objects based on example images. The system is outlined in Fig. 8 with an example of sorting Rubik’s cubes. First, the front RGB camera takes images of the example objects, which are then segmented with SAM [39] and built as class representative prototypes. Next, our DE-ViT detects objects and generates instance segmentation masks from the RGB image of a side-view RGBD camera. Then, the instance segmentation masks are combined with the depth image to produce the point-cloud of each object instance. The grasp pose for each object is generated with a heuristic-based pose generator [40]. We use a Kuka LBR iiwa robot. Note that our system receives example images of novel classes and detects novel objects instantly as it does not require finetuning.
Results. Tab. 5 shows the success rates of our system in object sorting. We adopt ten YCB objects such as crackers and mustard bottles. For each run, all objects are placed on the table, and we ask the robot to pick and place objects based on a given order of the novel classes. We use our DE-ViT with the ViT-L backbone. We use 3 example images for each object class, except the mustard bottle and tomato soup can, for which we use 6 images to improve their detection accuracy. Our system achieves an overall 97% success rate out of 100 independent picks for all objects. The layout of all objects is randomized for each pick. Of the three failed cases, two involve mistaking chips for a tomato can and vice versa. In the third case, the chips are detected correctly, but the instance mask is not precise enough to enable a successful grasp. We believe this is because our DE-ViT model is trained exclusively on the base categories of the LVIS dataset without using YCB images or any datasets on retail products.
4 Final Remarks
We conducted extensive analyses and ablation studies on model efficiency, comparisons to DINOv2 in Meta RCNN, comparisons to language-based detectors [32], and the effects of the number of PL layers and shots. Details of these studies and results are included in Appendix C due to space limit.
Appendix A Appendix
A.1 Code and Pretrained Models
The source code of our method DE-ViT is included in the supplementary folder Code. Please check Code/README.md for instructions on installation, dataset setup, and downloading pretrained models from an anonymous server.
Appendix B Related work
Few-shot Object Detection (FSOD) aims at detecting objects of novel classes by utilizing a few support images from novel classes as training samples [2]. Existing approaches can be broadly classified into finetuning-based [13, 25, 24, 27, 12] and meta-learning-based strategies [17, 22, 23]. Finetuning-based methods, despite their prevalence, suffer from a large accuracy gap between the base and novel classes, as well as practical limitations due to redundant multi-stage procedures [3]. Meta-learning-based methods avoid finetuning through online adaptation. Early works on meta-learning FSOD (Fig. 2.a) construct prototypes from novel class examples and use the prototypes as input to a network that classifies query images [17, 22]. Recent methods (Fig. 2.b) design interaction mechanisms of dense spatial features between query and support images [15, 16, 26]. In contrast, our method (Fig. 2.c) computes dot-products with the prototypes to project their features into a subspace, instead of using a prototype classifier, and applies a region-propagation network instead of a dense feature matching network. One-shot Object Detection (OSOD) is an extreme case of FSOD with only one example per novel class, it reduces the setting to single-class detection [11]. Prior approaches primarily focus on designing interaction mechanisms of dense spatial features between support and target images [30, 37, 41]. However, the OSOD formulation restricts the use of additional support and requires a separate inference per class [3, 38]. Compared with existing work, our method does not use per-class inference and only utilizes class-level prototypes without dense feature interactions.
Vision Transformer (ViT) is a recent architecture that demonstrates stronger performance and more interpretable features than traditional convolutional architectures. There is a growing trend of applying self-supervised ViTs, such as DINO [42, 33], to unsupervised object discovery [43, 44, 45, 5]. TokenCut [46] applies graph-cut over DINO features to separate the primary foreground object. DeepSpectral [4] predicts segmentation masks through unsupervised spectral clustering over DINO features. OW-DETR [47] uses DINO ViT to discover unknown objects in an open-world setting.
Robotics Application of Few-shot Object Detection. FSOD has been increasingly deployed in robotics applications [48]. AirInteraction [49] designs a robot exploration strategy based on human-informed interestingness, leveraging the learning from examples ability of few-shot detectors. CI-FSOD [50] designs a lightweight few-shot adaptor module that is suitable for robotics deployment on edge devices. Fewsol [51] studies the existing few-shot detector performance in typical robotics environments. TFOD [52] designs a mobile manipulator application purely based on the predictions of a few-shot detector with online annotations.
Appendix C Analysis and Ablation Studies
@@
Efficiency. We compare the inference time of DE-ViT with different values of $T$ against recent few-shot works in Tab. A1 in COCO 10-shots setting, where Swin denotes Swin Transformer [53], RN101 denotes ResNet101 [54], MFRCNN denotes Meta Faster RCNN [15], and CrossT denotes Cross Transformer [16]. $T$ is the number of classes to create subspace features and perform inference. Tab. A1 shows that our method DE-ViT can achieve the smallest inference time while having better accuracy. The inference times of all the compared methods are measured on the same machine.
@@
Table A1: Ablation study on different values of $T$ and inference time comparison with existing methods. N/A: The pretrained model is not publicly available for evaluation.
@@
Table A2: Comparison to naively applying DINOv2 with Meta RCNN, and ablation studies on the feature subspace projection.
{tabu}
to c|ccc|ccccMeta RCNN + Feature Subspace Projection   Novel Base
DINOv2 Backbone       $c$          $\mathcal{B}$    $\mathcal{C}\setminus c$ nAP50 nAP75 bAP50 bAP75
✓ 4.5 2.2 48.9 22.5
      ✓ 26.2 9.7 29.3 12
      ✓         ✓ 38.4 23 43.4 26.8
      ✓         ✓   ✓39.5 24.1 42.3 25.9
Table A3: Ablation studies on the region propagation-based localization.
Configuration Novel Base
nAP50 nAP75 bAP50 bAP75
Conventional regression network 37.7 14.6 46.5 23.9
Conventional regression network + expanded proposal 35.6 12.5 41.3 19.8
Region propagation network 39.5 24.1 42.3 25.9
Feature subspace projection. We examine the component effects for feature subspace projection in Tab. C. The first column represents the integration of raw features of DINOv2 ViT into Meta RCNN [17], a standard prototype-based few-shot detector. Tab. C shows that the model has an accuracy collapse on novel classes with raw ViT feature inputs by completely overfitting base classes. The $c$, $\mathcal{B}$, and $\mathcal{C}\setminus c$ represents different groups of prototypes used in feature projection, as in Eq. 5. By projecting features to $p_{c}$, general detection ability emerges (nAP50: 4.5 $\rightarrow$ 26.2), then significantly improves after adding background classes prototypes $p_{\mathcal{B}}$ (nAP50: 26.2 $\rightarrow$ 38.4), and further improves with $p_{\mathcal{C}\setminus c}$ (nAP50: 38.4 $\rightarrow$ 39.5). Results in Tab. C and A3 are obtained on COCO 2017 with ViT-S/14.
Region propagation network. We study the impacts of region-propagation-based localization in Tab. A3. The conventional regression network exhibits poor localization accuracy on novel classes (nAP75: 14.6). Note that a set of learned 2d positional embeddings is added to the ViT features before feeding into the conventional regression network. When using our proposed region-propagation network, the localization accuracy has a significant boost (nAP75: 14.6 $\rightarrow$ 24.1). Since our region-propagation network expands proposals, the accuracy gain may come from the spatial features of larger areas. Therefore, we ablate this by expanding proposals within the conventional regression network. This results in even lower performance than the conventional regression network alone (nAP75: 14.6 $\rightarrow$ 12.5). This shows our region propagation network is the main contributing factor to the performance boost. Note that the results in Tab. A3 are obtained by only changing the localization architectures while keeping the feature subspace projection.
Comparison to language-based detectors. Tab. A4 compares our method DE-ViT against language-based detectors on the COCO and LVIS-v1 datasets. Language-based detection, also known as open-vocabulary detection [55], does not assume access to support images on novel classes but requires the knowledge of the novel class names. The class names are then used to discover images of novel classes from a large collection of image-text pairs. Therefore, language-based detectors have the following limitations. First, many objects lack clear names [56], e.g., objects that are specific to certain contexts [57]. Second, the association between visual concepts and language is evolving and not static [58]. In contrast, few-shot object detection (FSOD) does not assume any knowledge of class names, and novel classes are described with support images only. FSOD’s setting is therefore more general and arguably more challenging, and aims to emulate humans’ capability to recognize objects by their consistent appearance.
Table A4: Comparison to language-based detectors on LVIS and COCO 2017. $\dagger$: use customized pretrained backbone instead of public ones. “-”: result that was not reported.
@@
DE-ViT (Ours) 5 shots ViT-S/14 20.6 19.6 28.3
ViT-B/14 26.4 25.0 36.1
ViT-L/14 28.5 27.9 38.3
10 shots ViT-S/14 21.8 21.9 36.1
ViT-B/14 29.9 28.1 42.4
ViT-L/14 32.4 31.9 46.2
30 shots ViT-S/14 24.2 23.4 39.5
@@
In COCO, our method DE-ViT outperforms the previous SoTA CORA+ by 6.9 AP50. Our method only trains on COCO while CORA+ uses ImageNet-21K [65] and COCO Captions [66] as additional training data. In LVIS, DE-ViT outperforms the previous SoTA on mask APr (+1.5 over F-VLM) and box APr (+2.4 over OWL-ViT). Note that we report the Co-Det [64] result with Swin-B backbone instead of EVA02-L backbone [67] because the latter includes the base and novel classes of COCO in its training set. Moreover, we observe a high variance in the performance of language-based detectors. F-VLM achieves a mask APr of 32.8 on LVIS but only has 28 nAP50 on COCO. While CORA+ has 43.1 nAP50 on COCO but only a box APr of 28.1 on LVIS. In contrast, our DE-ViT outperforms existing solutions on both LVIS and COCO. MM-OVOD (10-shots) [10] denotes the MM-OVOD language detector with 10 images per novel classes provided as examples.
@@
Background classes set $\mathcal{B}$. We study the impacts of using different background classes sets $\mathcal{B}$ in Tab. A6. The first and second rows represent using the stuff (non-object) classes, e.g., floor, sky, of COCOstuff [34] and ADE20k [68] to construct $\mathcal{B}$. Note that COCOstuff has 53 stuff classes and ADE20k has 35 stuff classes, which explains the minor accuracy drop from ADE20k. The third row denotes using all classes of ADE20k as $\mathcal{B}$. This indicates that adding thing (object) classes to $\mathcal{B}$ may not be beneficial. Tab. A6 shows that changing $\mathcal{B}$ only results in small accuracy variations. This suggests that our proposed feature subspace projection does not rely on specific prototypes.
Table A6: Ablation studies on the background classes set $\mathcal{B}$.
{tabu}
l|ccBackground Classes $\mathcal{B}$ nAP50 bAP50
COCOstuff [34] 43.1 45.6
ADE20k [68] 42.5 45.8
ADE20k (all) 42.1 44.7
More Shots. We study the model performance with different numbers of shots in Fig. 7 and Fig. A2, on COCO 2014 and COCO 2017, correspondingly. For COCO 2014, the numbers of shots are set from 1 to 10, 15, 20, and 30. To align with existing work, we use the same support images by previous work [13] for shots 2,3,5,10,30. For other shots, we sample within the support images mentioned above. For COCO 2017, we follow the conventional base/novel class splits of the one-shot benchmark. To measure with more robustness, we randomly select support images within the validation set of COCO 2017 for each query image, and compute nAP50 for all four novel class splits. The reported nAP50 is the average among all splits and choices of support images. The numbers of shots are set from 1 to 10, 15, 20, 30, 40, 50, 75, 100.
Table A7: Comparison of training epoch and parameter size to other detectors trained on LVIS.
Total
Params
Trained
Params
Epochs APr
OWL-ViT [36] 433M 433M 1800 31.2
F-VLM [69] 445M 25M 118 32.8
DE-ViT (Ours) 350M 23M 14.4 34.3
Figure A2: Detection accuracy under different numbers of shots in COCO 2017.
Using Boxes or Masks to Build Prototypes. We study the effects of annotation types such as bounding boxes or masks for prototype construction in Tab. LABEL:tab:ablation-bbox. We observe that using bounding boxes to build prototypes yields almost indistinguishable performance compared to using instance masks at even 5-shot in COCO.
Training Epochs and Parameter Sizes. In Tab. A7, we compare the parameter sizes and training epochs of detectors trained on the large-scale dataset LVIS. Tab. A7 shows that our method DE-ViT only has 23M trainable parameters, and is trained orders of magnitude faster than F-VLM [69] and OWL-ViT [36].
Hyperparameters. We report the hyperparameters used in our experiments in Tab. A8.
Table A8: Hyperparameters used in our experiments.
Hyperparameter Value
Max image size $800\times 1333\times 3$
Batch size 16
Learning rate 0.002
Optimizer SGD
Warmup steps 5000
Learning rate schedule stepwise decay
Learning rate decay steps 60000, 80000
Learning rate decay gamma 0.1
Training steps 90000
Evaluation frequency 5000 steps for COCO, 10000 steps for LVIS
Vision transformer size (layers, channel size) small (12, 384), base (12, 784), large (24, 1024)
Channels of propagation network 256
Channels of feature subspace projection ($S$) 256
Classes in feature subspace projection ($T$) 10
Number of PL layers 3 for COCO, 5 for LVIS
Preparation of YCB Objects. Fig. 1 shows the detection results of DE-ViT on YCB objects, a standard set of objects widely used in robotic manipulation benchmark [6]. There are misclassifications and inaccurate boxes, e.g., the white skillet is mistaken as a can, all round-shape fruits are recognized as orange, while the red one is an apple. However, we believe the overall result is encouraging. The specification of YCB objects at the time this paper is written includes 72 categories. We use a total of 33 by selecting and merging certain categories. The categories in use are apple, ball, banana, bowl, brick, can, cheez-it, chips, clamp, cleanser bottle, coffee jar, comet pine, cups, drill, glass, lego, lemon, marker, mug, mustard, orange, peach, pear, peg-hole, pitcher, plate, screwdriver, skillet, spray bottle, sugar box, toy airplane box, utensil, wood blocks jar. The source image in Fig. 1 is taken from the banner picture of ycbbenchmarks.com. For each category, we use Google Image Search to collect a few sample images. Fewer than four images on average are gathered per category. We annotate the corresponding objects by instance masks in each image using the software provided by SimpleClick [70]. Similar to SAM, SimpleClick generates instance masks automatically from user clicks, which significantly simplifies and accelerates the annotation procedure. Our annotator feedback indicates that annotating masks with SimpleClick is even easier and more accurate than drawing bounding boxes. An NVIDIA 3060 GPU is used for SimpleClick software. Class prototypes for YCB objects are built from the annotated example images. During DE-ViT inference, we replace prototypes of LVIS categories with those of YCB objects in order to detect these new categories. During postprocessing, We apply class-agnostic NMS and filter small bounding boxes. The data used for demonstration will be released upon acceptance.
Appendix D Limitations
In this work, we propose DE-ViT, a few-shot detector that uses example images to detect novel classes without any finetuning. We demonstrate that DE-ViT establishes new state-of-the-art in few-shot and one-shot benchmarks. One of the limitations is that our method DE-ViT uses a hybrid architecture of ViT and RCNN. The region proposal network (RPN) of RCNN is trained with only base classes and frozen during inference. This potentially limits the detection performance of novel objects. To overcome this limitation, one solution is to implement the RPN with our proposed framework as well. In doing so, the object proposals can be generated by conditioning on the prototypes of novel objects. Another solution is to adopt a full transformer architecture such as DETR [71] to eliminate the object proposal stage.
Our method DE-ViT adopts prototype-feature interactions, instead of dense spatial feature interactions that are computationally heavy. In doing so, our method DE-ViT archives greater accuracy and efficiency. However, class representative prototypes can potentially carry stronger bias from the pretrained ViTs such as DINOv2 than low-level spatial features. This would affect the model performance on long-tailed domains where the existing feature extractors underperform. A promising future direction is to design a more adaptive and bias-aware architecture to balance the efficient prototype-feature interactions and expensive spatial feature interactions.